Mira is five years older than Tanya. In three years, twice Mira's age will be triple Tanya's age. How old is Mira now?
Answer by stanbon(75887)

Mira is five years older than Tanya. In three years, twice Mira's age will be triple Tanya's age. How old is Mira now?
----
Equations:
m = t + 5
2(m+3) = 3(t+3)
--------
Substitute for "m" and solve for "t":
2(t+8) = 3t+9
2t+16 = 3t+9
t = 7 yrs
-----
Ans::
m = t+5
Mira = 7+5 = 12 yrs
--------------------
Cheers,
Stan H.
---------------